# Reinforcement Learning for Chain of Thought Compression with One-Domain-to-All Generalization

### v1

@@ Abstract
Chain-of-thought reasoning in large language models often creates an “overthinking trap,” leading to excessive computational cost and latency for unreliable accuracy gains. Prior work has typically relied on global, static controls that risk penalizing necessary reasoning. We introduce a sample-level, soft reinforcement learning compression method that penalizes inefficiently long rollouts, but only on problems where the model has already mastered and already produced a more concise rollout. Our experiments show that this method reduces average response length by 20–40% with comparable or higher accuracy. Crucially, the compression exhibits strong cross-domain generalization; a model trained on math spontaneously shortens responses on unseen tasks like code, instruction following, and general knowledge QA, with stable or improved accuracy. We demonstrate a stable post-training curriculum (accuracy-compression-accuracy) that can ultimately produce models that are more accurate and reason more concisely, arguing that such compression method should be a standard phase in developing efficient reasoning models.
@@ 1 Introduction
Generating long chains of thought (CoT) has significantly improved the complex reasoning abilities of large language models (weiChainofthoughtPromptingElicits2022; guoDeepSeekR1IncentivizesReasoning2025). However, this success introduces an “overthinking trap,” characterized by high computational cost, increased latency, and verbose reasoning paths that do not always guarantee higher accuracy (teamKimiK15Scaling2025; yuDAPOOpenSourceLLM2025). We posit that true intelligence manifests as compression. Once a model masters a problem-solving path, a more efficient representation must exist. We argue that compression is not merely a trade-off for efficiency but an indicator of internalized, mature capability.
Prevailing approaches, including reinforcement learning with various length penalties (teamKimiK15Scaling2025; yuDAPOOpenSourceLLM2025) and various forms of distillation (hsiehDistillingStepbyStepOutperforming2023; daiCaptureKeyReasoning2025; liTeachingSmallLanguage2024), have laid a crucial foundation. They demonstrate the feasibility of compressing CoT and offer valuable mechanisms for controlling computational budgets, often achieving effective accuracy-efficiency trade-offs. However, these methods typically rely on global and static controls, such as a uniform penalty or a fixed truncation limit.
@@
Our experiments reveal several key findings. First, this sample-level, gated, soft compression framework significantly outperforms global penalties, while hard truncation methods fail catastrophically. The gating mechanism — compressing only after mastery is achieved — proves essential for maintaining robust performance. Second, this compression skill generalizes: models trained to compress reasoning on a specific domain (e.g., mathematics) spontaneously reduce verbosity across all tasks. Third, we identify a stable, multi-stage training recipe which can reduce average inference length by 20–40% with comparable or even improved accuracy, a robust result across model structures, scales, and tasks.
@@
Our contributions are threefold. We (1) propose a minimal, stable, and sample-level gated compression paradigm, formalized as a practical post-training recipe; (2) systematically validate the mechanisms, from motivation to empirical phenomena, providing reproducible evidence of its dynamics, generalization, and failure modes, leading to a robust early stopping strategy; (3) reframe compression from a mere cost-optimization technique to a measurable process of capability internalization, delivering a standardized training phase and a tunable control for model efficiency.
@@ 2 Related Work
Most prior work relies on global, static policies-hard truncation (ThinkPrune (houThinkPrunePruningLong2025)), truncation-penalized RL (DLER (liuDLERDoingLength2025)), length-harmonizing fine-tuning (O1-Pruner (luoO1PrunerLengthHarmonizingFineTuning2025)), constraint-following RL (L1 (aggarwalL1ControllingHow2025)), sample-level non-gated length penalties RL in Kimi K1.5 (teamKimiK15Scaling2025), and global non-gated length penalties in DAPO (yuDAPOOpenSourceLLM2025) — whereas we apply sample-level, gated, soft compression that triggers only after mastery and only beyond a self-derived safe length.
@@
CoT distillation (Distilling Step-by-Step (hsiehDistillingStepbyStepOutperforming2023), EDIT (daiCaptureKeyReasoning2025), D&R Distillation (liTeachingSmallLanguage2024)), token/computation pruning and compression (Efficient Long CoT (wangEfficientLongCoT2025), DLCoT (luoDeconstructingLongChainofThought2025), TokenSkip (xiaTokenSkipControllableChainofThought2025), LightThinker (zhangLightThinkerThinkingStepbyStep2025), Prune-on-Logic (zhaoCanPruningImprove2025)), and model merging (Thinking Spectrum (lanThinkingSpectrumEmpirical2025), model merging for length-controlled reasoning (wuUnlockingEfficientLongtoShort2025)) pursue similar accuracy-efficiency goals, yet our loop is on-policy and compresses only when the model has demonstrably mastered the instance.
@@ 3 Method
We adopt the standard reinforcement learning with verifiable rewards (RLVR) training strategy (guoDeepSeekR1IncentivizesReasoning2025). We compress only where the model already knows the answer. For each training sample, tokens generated beyond a per-sample “safe length” are softly penalized, but only after the model has demonstrated reliable mastery in solving that sample.
@@
We then apply a soft penalty $r^{\text{len}}(x,y)$ as a piecewise linear function similar to DAPO (yuDAPOOpenSourceLLM2025). Responses shorter than the safe median length ($L^{\text{start}}$) are not penalized. Responses longer than the maximum observed correct length ($L^{\text{max}}$) receive a full penalty of 1-1. Responses with lengths between these two bounds are penalized linearly:
@@
We optimize the model using the GRPO algorithm family (including GRPO (shaoDeepSeekMathPushingLimits2024), DAPO (yuDAPOOpenSourceLLM2025), Dr. GRPO (liuUnderstandingR1ZeroLikeTraining2025), and modifications for mixture-of-experts models (zhengGroupSequencePolicy2025; maStabilizingMoEReinforcement2025)) of policy-gradient methods, which are robust for our setting. Additionally, we apply early stopping during the compression phase: we monitor validation accuracy and average response length, stopping at the “pre-collapse” optimum — the point where accuracy first shows a sustained decline while length is still falling. This reliably avoids the over-compression regime.
@@ 4.1 Experimental Setup
We base our experiments on MiMo-7B-RL (xiaomiMiMoUnlockingReasoning2025), a 7B dense thinking model with strong mathematical reasoning capabilities but relatively long CoTs. During the training, we set its context length to 64K tokens to accommodate long-form mathematical reasoning. We use 113K math problems with ground truth answers for RLVR training, sourced from various public datasets.
Our training algorithm is DAPO (yuDAPOOpenSourceLLM2025), a GRPO-family policy gradient method, with clipping parameters $\text{clip\_low}=0.2$ and $\text{clip\_high}=0.27$. We implement a dynamic sampling and mixing strategy for training: each batch is constructed so that 90% of samples are used for standard capability training (pursuing correctness). The remaining 10% of samples, drawn from those the model has already mastered ($\text{passrate}=1.0$), are used for compression training via our sample-level soft penalty. Unlike DAPO, we do not filter out samples with passrates 0 or $1$ when constructing the training set. The batch size is set to 256 and the rollout size is $n=8$. We use AdamW optimizer (loshchilovDecoupledWeightDecay2018) for training. The learning rate is set to $3\times 10^{-6}$ with a $10$-step warmup. We use a training temperature of $1.0$ and a validation temperature of $0.6$, with top-$p$ set to $1.0$ for both.
We compare our method against two baselines:
@@
We evaluate performance on the AIME24 (MathaiAime24Datasets) and AIME25 (MathaiAime25Datasets) benchmarks, reporting correctness and average response length, with all metrics computed over 32 samples per test case and averaged (avg@32).
@@ 5 Detailed Analysis
  • First, we use a MiMo-7B-SFT model (xiaomiMiMoUnlockingReasoning2025) and apply our compression training using Dr.GPRO (liuUnderstandingR1ZeroLikeTraining2025), mixing varying ratios (10%, 20%, 40%, and 90%) of mastered samples ($\hat{p}(x)=1.0$) and remaining samples with $\hat{p}(x)\in(0.0,1.0)$ during the compression phase.
  • Second, to assess architectural generalization, we experiment with Qwen3-30B-A3B (yangQwen3TechnicalReport2025), a mixture-of-experts (MoE) model. This model is first doing supervised fine-tuning on open-source CoT data to inject the reasoning pattern, then trained in an accuracy-focused phase for 150 steps using Dr.GRPO with a rollout routing replay mechanism (maStabilizingMoEReinforcement2025). Then, we undergo two full loops of the accuracy $\rightarrow$ compression curriculum, again with a 40% compression mixture.
@@
A key observation across models is the effect of optimizer (e.g., from AdamW (loshchilovDecoupledWeightDecay2018)). After switching from compression back to accuracy, the model briefly continues to shorten responses and may see a transient accuracy drop before the optimization objective stabilizes. For instance, the MiMo-7B model required 15 steps post-compression before accuracy began to recover and 25 steps before the length started to increase.
@@
AIME25 (MathaiAime25Datasets) Math 1.7-1.7 31.7-31.7
MATH500 (HuggingFaceH4MATH500Datasets2025) Math 0.8-0.8 41.6-41.6
GPQADiamond (FingertapGPQADiamondDatasets) QA 0.4-0.4 34.8-34.8
MMLU_Redux (gemaAreWeDone2025) General Knowledge +1.1+1.1 33.4-33.4
IFEval (zhouInstructionFollowingEvaluationLarge2023) Instruction Following +0.9+0.9 22.8-22.8
LiveCodeBenchv6 (jainLiveCodeBenchHolisticContamination2024) Code +3.7+3.7 31.8-31.8
ZebraLogic (linZebraLogicScalingLimits2025) Logic 1.3-1.3 22.1-22.1
@@
We return, in the very last, to our opening claim that intelligence manifests as compression. Borges writes in Funes the Memorious, “To think is to forget differences, to generalize, to abstract.” In our mastery-gated RL objective, the model learns to reason at length, then to forget what does not matter, so that compressed chains of thought on mastered math tasks become abstractions that transfer across domains. Generalization, in this sense, is just forgetting the right differences well.

### v2

@@ Abstract
Chain-of-thought reasoning in large language models can trigger an “overthinking trap”: longer rollouts raise cost and latency yet often yield unreliable accuracy gains. Existing methods use global, static controls that may suppress needed reasoning. We propose mastery-gated, sample-level, soft reinforcement learning compression that penalizes long rollouts only when the model already solves the problem and has produced a shorter rollout. Across benchmarks, it cuts response length by 20-40% with comparable or higher accuracy and generalizes across domains: a model trained on math spontaneously shortens unseen tasks (code, instruction following, general-knowledge QA) without hurting accuracy. We further show two-way transfer between non-agent CoT and tool-use agents: non-agent training reduces SWE-Bench Verified rounds by 13%, while compressing a thinking agent cuts SWE trajectories by 67% tokens and 52% rounds and shortens non-agent outputs by up to 44%. Compression is thus not cosmetic brevity, but an inherent computation policy — what to keep, and what to forget.
@@ 1 Introduction
Generating long chains of thought (CoT) has significantly improved the complex reasoning abilities of large language models [weiChainofthoughtPromptingElicits2022, guoDeepSeekR1IncentivizesReasoning2025]. However, this success introduces an “overthinking trap,” characterized by high computational cost, increased latency, and verbose reasoning paths that do not always guarantee higher accuracy [teamKimiK15Scaling2025, yuDAPOOpenSourceLLM2025]. We posit that true intelligence manifests as compression. Once a model masters a problem-solving path, a more efficient representation must exist. We argue that compression is not merely a trade-off for efficiency but an indicator of internalized, mature capability.
Prevailing approaches, including reinforcement learning with various length penalties [teamKimiK15Scaling2025, yuDAPOOpenSourceLLM2025] and various forms of distillation [hsiehDistillingStepbyStepOutperforming2023, daiCaptureKeyReasoning2025, liTeachingSmallLanguage2024], have laid a crucial foundation. They demonstrate the feasibility of compressing CoT and offer valuable mechanisms for controlling computational budgets, often achieving effective accuracy-efficiency trade-offs. However, these methods typically rely on global and static controls, such as a uniform penalty or a fixed truncation limit.
@@
Our experiments reveal several key findings. First, this sample-level, gated, soft compression framework significantly outperforms global penalties, while hard truncation methods fail catastrophically. The gating mechanism — compressing only after mastery is achieved — proves essential for maintaining robust performance. Second, this compression skill generalizes: models trained to compress reasoning on a specific domain (e.g., mathematics) spontaneously reduce verbosity across all tasks. Third, we identify a stable, multi-stage training recipe which can reduce average inference length by 20–40% with comparable or even improved accuracy, a robust result across model structures, scales, and tasks. Finally, we show that this learned notion of “efficient computation” extends beyond single-turn CoT: it transfers to tool-use agents and even exhibits two-way generalization between non-agent reasoning and agent trajectories.
@@
Our contributions are threefold. We (1) propose a minimal, stable, and sample-level gated compression paradigm, formalized as a practical post-training recipe; (2) systematically validate the mechanisms, from motivation to empirical phenomena, providing reproducible evidence of its dynamics, generalization, and failure modes, leading to a robust early stopping strategy; (3) reframe compression from a mere cost-optimization technique to a measurable process of capability internalization, and show it as an agent-agnostic efficiency skill that applies to both pure reasoning and tool-use settings.
@@ 2 Related Work
Most prior work relies on global, static policies-hard truncation (ThinkPrune [houThinkPrunePruningLong2025]), truncation-penalized RL (DLER [liuDLERDoingLength2025]), length-harmonizing fine-tuning (O1-Pruner [luoO1PrunerLengthHarmonizingFineTuning2025]), constraint-following RL (L1 [aggarwalL1ControllingHow2025]), sample-level non-gated length penalties RL in Kimi K1.5 [teamKimiK15Scaling2025], and global non-gated length penalties in DAPO [yuDAPOOpenSourceLLM2025] — whereas we apply sample-level, gated, soft compression that triggers only after mastery and only beyond a self-derived safe length.
@@
CoT distillation (Distilling Step-by-Step [hsiehDistillingStepbyStepOutperforming2023], EDIT [daiCaptureKeyReasoning2025], D&R Distillation [liTeachingSmallLanguage2024]), token/computation pruning and compression (Efficient Long CoT [wangEfficientLongCoT2025], DLCoT [luoDeconstructingLongChainofThought2025], TokenSkip [xiaTokenSkipControllableChainofThought2025], LightThinker [zhangLightThinkerThinkingStepbyStep2025], Prune-on-Logic [zhaoCanPruningImprove2025]), and model merging (Thinking Spectrum [lanThinkingSpectrumEmpirical2025], model merging for length-controlled reasoning [wuUnlockingEfficientLongtoShort2025]) pursue similar accuracy-efficiency goals, yet our loop is on-policy and compresses only when the model has demonstrably mastered the instance.
@@ 3 Method
We adopt the standard reinforcement learning with verifiable rewards (RLVR) training strategy [guoDeepSeekR1IncentivizesReasoning2025]. We compress only where the model already knows the answer. For each training sample, tokens generated beyond a per-sample “safe length” are softly penalized, but only after the model has demonstrated reliable mastery in solving that sample.
@@
We then apply a soft penalty $r^{\text{len}}(x,y)$ as a piecewise linear function similar to DAPO [yuDAPOOpenSourceLLM2025]. Responses shorter than the safe median length ($L^{\text{start}}$) are not penalized. Responses longer than the maximum observed correct length ($L^{\text{max}}$) receive a full penalty of 1-1. Responses with lengths between these two bounds are penalized linearly:
@@
We optimize the model using the GRPO algorithm family (including GRPO [shaoDeepSeekMathPushingLimits2024], DAPO [yuDAPOOpenSourceLLM2025], Dr. GRPO [liuUnderstandingR1ZeroLikeTraining2025], and modifications for mixture-of-experts models [zhengGroupSequencePolicy2025, maStabilizingMoEReinforcement2025]) of policy-gradient methods, which are robust for our setting. Additionally, we apply early stopping during the compression phase: we monitor validation accuracy and average response length, stopping at the “pre-collapse” optimum — the point where accuracy first shows a sustained decline while length is still falling. This reliably avoids the over-compression regime.
@@ 4.1 Experimental Setup
We base our experiments on MiMo-7B-RL [xiaomiMiMoUnlockingReasoning2025], a 7B dense thinking model with strong mathematical reasoning capabilities but relatively long CoTs. During the training, we set its context length to 64K tokens to accommodate long-form mathematical reasoning. We use 113K math problems with ground truth answers for RLVR training, sourced from various public datasets.
Our training algorithm is DAPO [yuDAPOOpenSourceLLM2025], a GRPO-family policy gradient method, with clipping parameters $\text{clip\_low}=0.2$ and $\text{clip\_high}=0.27$. We implement a dynamic sampling and mixing strategy for training: each batch is constructed so that 90% of samples are used for standard capability training (pursuing correctness). The remaining 10% of samples, drawn from those the model has already mastered ($\text{passrate}=1.0$), are used for compression training via our sample-level soft penalty. Unlike DAPO, we do not filter out samples with passrates 0 or $1$ when constructing the training set. The batch size is set to 256 and the rollout size is $n=8$. We use AdamW optimizer [loshchilovDecoupledWeightDecay2018] for training. The learning rate is set to $3\times 10^{-6}$ with a $10$-step warmup. We use a training temperature of $1.0$ and a validation temperature of $0.6$, with top-$p$ set to $1.0$ for both.
We compare our method against three baselines:
@@
We evaluate performance on the AIME24 [MathaiAime24Datasets] and AIME25 [MathaiAime25Datasets] benchmarks, reporting correctness and average response length, with all metrics computed over 32 samples per test case and averaged (avg@32).
@@ 5 Detailed Analysis
  • First, we use a MiMo-7B-SFT model [xiaomiMiMoUnlockingReasoning2025] and apply our compression training using Dr.GPRO [liuUnderstandingR1ZeroLikeTraining2025], mixing varying ratios (10%, 20%, 40%, and 90%) of mastered samples ($\hat{p}(x)=1.0$) and remaining samples with $\hat{p}(x)\in(0.0,1.0)$ during the compression phase.
  • Second, to assess architectural generalization, we experiment with Qwen3-30B-A3B [yangQwen3TechnicalReport2025], a mixture-of-experts (MoE) model. This model is first doing supervised fine-tuning on open-source CoT data to inject the reasoning pattern, then trained in an accuracy-focused phase for 150 steps using Dr.GRPO with a rollout routing replay mechanism [maStabilizingMoEReinforcement2025]. Then, we undergo two full loops of the accuracy $\rightarrow$ compression curriculum, again with a 40% compression mixture.
@@
A key observation across models is the effect of optimizer (e.g., from AdamW [loshchilovDecoupledWeightDecay2018]). After switching from compression back to accuracy, the model briefly continues to shorten responses and may see a transient accuracy drop before the optimization objective stabilizes. For instance, the MiMo-7B model required 15 steps post-compression before accuracy began to recover and 25 steps before the length started to increase.
@@
AIME25 [MathaiAime25Datasets] Math 1.7-1.7 31.7-31.7
MATH500 [HuggingFaceH4MATH500Datasets2025] Math 0.8-0.8 41.6-41.6
GPQADiamond [FingertapGPQADiamondDatasets] QA 0.4-0.4 34.8-34.8
MMLU_Redux [gemaAreWeDone2025] General Knowledge +1.1+1.1 33.4-33.4
IFEval [zhouInstructionFollowingEvaluationLarge2023] Instruction Following +0.9+0.9 22.8-22.8
LiveCodeBenchv6 [jainLiveCodeBenchHolisticContamination2024] Code +3.7+3.7 31.8-31.8
ZebraLogic [linZebraLogicScalingLimits2025] Logic 1.3-1.3 22.1-22.1
@@
6 From CoT Compression to Agent Compression
Our method is motivated by compressing chain-of-thought (CoT) without sacrificing correctness. This framing is not specific to single-shot CoT. In this section, we show that the same recipe directly extends to agentic settings, where the model must interleave reasoning with tool use (e.g., edit–test loops), and that compression exhibits strong two-way generalization between non-agent and agent models.
Setup
We use MiMo-V2-Flash [teamMiMoV2FlashTechnicalReport2026] as the base model (309B total, 15B active parameters) and run two compression experiments. (i) Non-agent compression: we train the compression stage on non-agent math, code, and logic tasks for 80 steps. (ii) Agent compression: starting from a code-agent model optimized for thinking-mode via SFT and RLVR, we train the compression stage on thinking code-agent tasks for 100 steps with tool use enabled. In both cases, we apply the same mastery-gated RL objective from Section˜3, penalizing response length (not tool calls or rounds). We evaluate on three non-agent benchmarks (AIME25, LiveCodeBench-v6, GPQA-Diamond) and one agent benchmark (SWE-Bench Verified [openaiIntroducingSWEbenchVerified2024]).
Results
We observe strong two-way generalization between non-agent and agent compression.
Non-agent to agent (round reduction without agent training). Evaluated on SWE-Bench Verified in non-thinking mode, even compression trained purely on non-agent tasks induces more concise agent behavior: the non-thinking SWE agent reduces its average rounds from 110 to 96 ($\Delta R=-13\%$) with only a small accuracy change (73.4 $\rightarrow$ 72.3). This suggests that the learned compression behavior is not confined to single-shot CoT formatting, but can alter the agent’s interaction policy.
Agent compression (shorter trajectories and cross-domain transfer). When we compress the thinking-mode agent directly, we obtain large reductions in both trajectory length and rounds on SWE-Bench Verified: it reduces length from 105k to 34.2k tokens ($\Delta L=-67\%$) and rounds from 96 to 46 ($\Delta R=-52\%$), with a modest accuracy drop (72.1 $\rightarrow$ 70.3). Beyond SWE, the same compressed agent also produces shorter responses on a diverse set of non-agent benchmarks, indicating strong transfer of the learned computation-compression behavior (Table˜3).
Benchmark base step-100
AIME25 86.5/29.3k 85.3/21.9k  ( $\Delta L=-25\%$ )
LiveCodeBench-AA 76.2/23.3k 76.0/17.7k  ( $\Delta L=-24\%$ )
GPQA-Diamond 83.3/24.9k 84.5/14.0k  ( $\Delta L=-44\%$ )
SWE-Bench Verified 72.1/105.0k/96 70.3/34.2k/46  ( $\Delta L=-67\%$, $\Delta R=-52\%$ )
Table 3: Agent compression: accuracy/response-length on non-agent benchmarks and accuracy/response-length/rounds on SWE-Bench Verified in thinking mode. Relative changes ($\Delta L$, $\Delta R$) are computed against the base checkpoint.
We return, at the very last, to our opening claim that intelligence manifests as compression. Borges writes in Funes the Memorious, “To think is to forget differences, to generalize, to abstract.” In our mastery-gated objective, the model learns first to spend computation where it is needed, and then to forget what does not matter: not only to compress chains of thought on mastered problems, but to compress the entire reasoning process—tokens, turns, and tool-use trajectories—into a sharper policy of action. On this view, compression is not a cosmetic shortening of explanations; it is a learned allocation of attention and effort. And generalization is the same act, lifted across tasks: preserving the invariants, discarding the incidental, and forgetting the right differences well.